OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5?

OpenStudy (anonymous):

y-1=5(x-4)

OpenStudy (lalaly):

y-1=5(x-4)

OpenStudy (anonymous):

y=5x-19

OpenStudy (akshay_budhkar):

y-1=5(x-4) use the slope point form

OpenStudy (akshay_budhkar):

oh i am late!!!